OpenStudy (anonymous):

What is the slope intercept form of the equation 3x-y= - 6? What is the slope intercept form of the equation x-3y=3? What is the slope intercept form of the equation 9x+3y=6? What is the slope intercept form of the equation x+3y=-6

OpenStudy (anonymous):

a = y = -1 ----x - 2 3 b= y = 1 --- x - 1 3 c = y - 3x+2 d = y -3x + 6

OpenStudy (anonymous):

recheck a and d.... b,c are right!!

OpenStudy (anonymous):

its basically solving for y...

OpenStudy (anonymous):

they are right

OpenStudy (anonymous):

a) y=3x+6 d) y=(1/3)x+2

OpenStudy (anonymous):

do you understand ur mistake???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you manipulate the equation in such a way that, y is on the Left hand side and the the others on the right hand side..... a) 3x-y=-6 -y=-3x-6 y=3x+6 d) x+3y=-6 3y=-x-6 y=-(1/3)x-6 \[y=-\frac{1}{3}x-6\]

OpenStudy (anonymous):

alright